OpenStudy (anonymous):

Jack's age next year will be twice Jill's age last year. Their present age will present ages total 45. How old is each now?

OpenStudy (anonymous):

... This is a duplicate...

OpenStudy (anonymous):

2 independent eqns and 2 unknowns. get x and y.

OpenStudy (anonymous):

@heyyygirrrl you got this?

OpenStudy (anonymous):

it is actually not that easy

OpenStudy (anonymous):

here is an easier way put jack's age as \(x\) and since the total is 45 jill's age must be \(45-x\)

OpenStudy (anonymous):

jack's age next year will be \(x+1\) and twice jills age last year will be \(2(44-x)\)

OpenStudy (anonymous):

solve \[x+1=2(44-x)\] in a couple steps \[x+1=88-2x\] \[3x+1=88\] \[3x=87\] and finally divide by \(3\)

OpenStudy (anonymous):

29